# Adaptive Informed Deep Neural Networks for Power Flow Analysis

### v1

@@ IV Proposed architecture: PINN4PF
Figure 1: Schematic diagram of PINN4PF. The architecture includes two advancements: (A) a double-head feed-forward NN with an adaptive activation function, and (B) a physics-based loss function.
@@ IV-A Forward Pass
A double-head feed-forward NN $f(\cdot)\in\{(f_{0},f_{1},f_{2})\}$ is developed to approximate $\hat{\vec{y}}=\{(\vec{\hat{\mu}}_{i},\vec{\hat{\omega}}_{i}):i=1,2,\dots,n\}$ at load buses using the dataset $\{\vec{x},\vec{y}\}$. It has three types of layers, i.e., input, hidden, and output layers, as shown in Fig. 1-A. The input and output layers correspond to the input features $\vec{x}$ and output labels $\vec{y}$, respectively. The neurons of the input layer contain active power $\vec{p}^{d}$ and reactive power $\vec{q}^{d}$ at all load buses, and hence, the input layer has $n\times 2$ neurons, where $n$ is the number of load buses. Following the input layer, there is a set of shared hidden layers $f_{0}(\cdot)$ acting as a feature extractor that projects $\vec{x}$ to a higher-dimensional space where the two heads, $f_{1}(\cdot)$ and $f_{2}(\cdot)$, separately involve a few hidden layers to respectively approximate $\vec{\mu}$ and $\vec{\omega}$. Thus, the output layer of each head has $n$ neurons. The shared set of hidden layers and the two heads each are a chain of functions and can be represented as
$f_{0}(\vec{x})=l_{m}^{0}\circ\ldots\circ l_{1}^{0}(\vec{x}),$ (5)
$f_{1}(\vec{x})=l_{m}^{1}\circ\ldots\circ l_{1}^{1}(\vec{x}),$ (6)
$f_{2}(\vec{x})=l_{m}^{2}\circ\ldots\circ l_{1}^{2}(\vec{x}),$ (7)
which sequentially process $\vec{x}$ through $m$ hidden layers to obtain $\vec{f}_{0}(\cdot)$, $\vec{f}_{1}(\cdot)$, and $\vec{f}_{2}(\cdot)$, respectively. For the shared set of hidden layers and the two heads, the $k$-th hidden layer ($k=1,2,\dots,m$) is given by
@@
The Rectified Linear Unit (ReLU) activation function $\sigma(\cdot)$ is applied to the shared set of hidden layers of the two heads (5). While an adaptive version of ReLU $\sigma^{\prime}(\cdot)$ is developed to apply model-based nonlinear transformations to the two heads (6)–(7). That is, a trainable parameter $\alpha$ is introduced to scale the result of the linear transformation while applying nonlinearity. For $\sigma^{\prime}(\cdot)$, the trainable parameter $\alpha$ serves as a guide to prevent overfitting and improve the generalization ability of the NN against unseen data. $\sigma(\cdot)$ and $\sigma^{\prime}(\cdot)$ are represented as
$\sigma=\max(0,z),$ (9)
$\sigma^{\prime}=\max(0,\alpha z),$ (10)
where $z$ is equivalent to $W_{k}^{T}\cdot\vec{x}+b_{k}$. During each iteration of the training process, the gradients of the loss function concerning the trainable parameters $\nabla_{W_{k},b_{k},\alpha_{k}}L$ are computed for each hidden layer $k$ using the chain rule of differentiation. The resulting gradients are then backpropagated through the network to update and optimize the weights $W_{k}$, biases $b_{k}$, and the scaling parameter $\alpha_{k}$. This process enables the network to learn the optimal values of $\alpha_{k}$, along with $W_{k}$ and $b_{k}$, potentially leading to improved performance in approximating $\hat{\vec{y}}$. More detailed information about adaptive activation functions can be found in [37, 38].
@@ IV-B Loss Function
The derivation starts with Ohm’s law, which relates voltage $V$, current $I$, and resistance $R$, that is, $V=I\times R$. By extending this equation to the complex domain, $\overline{V}=\overline{I}\times\overline{Z}$ relates the voltage phasor, current phasor, and impedance phasor. From this equation, $\overline{I}$ can be expressed in terms of the voltage phasor and admittance phasor $\overline{Y}$ (inverse of the impedance phasor), that is $\overline{I}=\overline{Y}\times\overline{V}$. The system of equations for all the buses can then be presented in matrix form as
@@
By rearranging (11), the current flowing into bus $k$ can be presented as a linear combination of the voltages at all other buses with weights given by the corresponding admittance matrix
@@
Here, the power equation $S=V\times I^{*}$ provides further guides for subsequent derivations that is $I={S^{*}\over{V^{*}}}$, where $S$ is complex apparent power. The current and voltage phasors are related as
@@
wherein the right-hand side is an extended form of the right-hand expression in (12). Finally, the expression for calculating the voltage phasor at bus $k$ can be written as
@@
In (14), $Y_{kk}=G_{kk}+jB_{kk}$ is known from the power system topology. Considering the input features and output labels needed to develop the NN for PF analysis, $V_{k}=\mu_{k}+j\omega_{k}$ and $V_{k}^{*}=\mu_{k}-j\omega_{k}$ are known from the output labels, and $S_{k}^{*}=p_{k}^{d}+jq_{k}^{d}$ is known from the input features. The only unknown expression remaining in (14) is
@@
which is called hidden function in this study. Note that $\psi$ is unique for each data point. Theoretically, $\psi$ is also unique for each bus. During the training process, the NN first approximates $\psi$ and then uses it to approximate $V=\mu+j\omega$ using
@@
This unique relation improves the learning process by following the underlying physical laws of the power system. Note that instead of integrating the whole admittance matrix $\begin{bmatrix}Y\end{bmatrix}_{n\times n}$, only the diagonal elements $Y_{kk}$ are needed to obtain the physical loss term, and therefore, the computational resource required is reduced by up to three orders of magnitude compared to when integrating the whole topology information.
Typically, the goal of the training process is to minimize the difference between $\hat{\vec{y}}$, approximated by the NN, and ground-truth output labels $\vec{y}$, obtained from the Newton-Raphson method (NR), based on a loss function of choice. Here, a modified loss function is developed that integrates both supervised and physical penalty terms. The former is the mean square of the difference between the output approximated by the NN $f(\cdot)$ and $\vec{y}$. While the latter is the mean square of the difference between the output obtained from the physical model $f^{\prime}(\cdot)$ and $\vec{y}$
@@
where $N$ is the total number of data points $j$. $f(\cdot)\in\{(f_{0},f_{1},f_{2})\}$ and $f^{\prime}(\cdot)$ are the NN architecture (5)–(7) and the physical model (16), respectively. $\beta_{0}$ and $\beta_{1}$ represent the coefficients for the supervised penalty term and the physical penalty term, respectively, with the constraint that $\beta_{0}+\beta_{1}=1$. Initially, $\beta_{0}=1$ and $\beta_{1}=0$. After 100 epochs, the value of $\beta_{1}$ increases at a specified rate, with its maximum value determined from the sensitivity analysis. $\theta=\{W,b\}$ is the trainable parameters of the NN and $\alpha$ is the trainable parameter for the activation function.
@@ V Results
The performance of PINN4PF is evaluated against two baselines: a linear regression model (LR) and a black-box NN (MLP). Experiments are performed on 4-bus [32], 15-bus [33], 290-bus [34], and 2224-bus [35] test systems. The selected systems contain one reference bus with known voltages in complex number form, $\mu_{0}+j\omega_{0}$, and unknown $p_{0}^{d}$ and $q_{0}^{d}$, and load buses $i$ for which $p_{i}^{d}$ and $q_{i}^{d}$ are known, while voltages $\mu_{i}+j\omega_{i}$ are unknown.
Input features and output labels of the datasets are represented as $\vec{x}$ and $\vec{y}$, respectively, where $\vec{x}$ includes $(\vec{p}^{d}_{i},\vec{q}^{d}_{i})$ and $\vec{y}$ includes $(\vec{\mu}_{i},\vec{\omega}_{i})$ obtained from the Newton-Raphson numerical method (NR) for all load buses $i$. The PandaPower Python package [39] is used to perform NR and generate the datasets. Note that PandaPower specifies the state variables of power systems, i.e., $[\delta\text{ }v]^{T}$, that is, there is a need for converting the state variables to $\mu_{i}=v_{i}cos\delta_{i}$ and $\omega_{i}=v_{i}sin\delta_{i}$ to yield the output labels $\vec{y}=\{(\vec{\mu}_{i},\vec{\omega}_{i}):i=1,2,\dots,n\}$.
@@ V-A Model Setup
The training process ends after 5000 epochs for PINN4PF, MLP, and LR. The loss function used is (17), with $\beta_{0}=1$ and $\beta_{1}=0$ for MLP and LR, indicating a supervised penalty term. However, $\beta_{0}$ and $\beta_{1}$ are non-zero for PINN4PF where $\beta_{0}+\beta_{1}=1$, indicating a combination of supervised and physical penalty terms. The activation function is ReLU (9) for MLP and the shared hidden layers of PINN4PF. The adaptive ReLU (10) is used for the separated hidden layers of PINN4PF. The Adam optimization algorithm updates the trainable parameters during the training process for PINN4PF, MLP, and LR.
@@ V-B2 Robustness
The robustness of PINN4PF, MLP, and LR is evaluated using the 15-bus test system under varying noise levels. Controlled noise levels ranging from 0% to 10% are introduced to both input features and output labels of the training dataset. The corrupted vectors are defined as $\vec{x}^{\prime}=\vec{x}\pm\vec{r}_{x}$ and $\vec{y}^{\prime}=\vec{y}\pm\vec{r}_{y}$, where $\vec{r}_{x}$ and $\vec{r}_{y}$ are vectors of random values between 0 and 1, and 0 and 0.1, respectively. Fig. 3 illustrates the changes in the MSE for the voltage magnitude $[V^{2}]$ based on the testing dataset with varying noise levels for MLP. The comparison is made relative to a constant curve, representing the MSE obtained for PINN4PF using the highest noise level, i.e., 10%. At the 0% noise level, MLP outperforms PINN4PF trained with 10% noisy data by approximately 18%. However, as the noise level increases, the performance of MLP performance deteriorates significantly, with MSE increasing up to six times. LR is excluded from the graph as its MSE at different noise levels is two orders of magnitude higher than that of PINN4PF. In addition, LR shows limited improvement with increasing dataset sizes and consistently underperforms compared to both PINN4PF and MLP.
@@ V-B3 Training dataset size
The impact of the size of the training dataset on the performance of MLP is investigated using the 15-bus test system. Fig. 4 illustrates the changes in the MSE for the voltage magnitude $[V^{2}]$ based on the testing dataset with varying training dataset sizes. The comparison is made relative to a constant curve representing the MSE obtained for PINN4PF using 256 training data points. It is observed that MLP requires a training dataset twice as large as that of PINN4PF to achieve a still inferior performance. However, the performance of MLP improves with more training data, reaching a comparable level with a dataset four times larger than that of PINN4PF. This indicates that PINN4PF is more data-efficient. LR is not included in the graph as its MSE with different training dataset sizes is two orders of magnitude larger than that of PINN4PF. Additionally, LR shows limited improvement with increasing training dataset sizes and consistently performs worse than PINN4PF and MLP. This indicates that the capacity of LR to capture complex relationships in PF analysis is significantly lower, and increasing the amount of training data is insufficient to bridge the performance gap.
@@ V-B4 Accuracy of derived power flow quantities
For derived physical quantities, i.e., line current, line active power, and line reactive power, the mean and standard deviation of the testing MSE obtained by PINN4PF and MLP are computed and compared in Fig. 5. It is also observed that PINN4PF achieves up to 81%, 63%, and 66% lower maximum testing MSE for the line current $[A^{2}]$, line active power $[W^{2}]$, and line reactive power $[VAR^{2}]$, respectively, compared to the MLP.
@@ V-B5 Scalability
The experiments are extended to compare the performance of PINN4PF, MLP, and LR across different test system sizes: 4-bus, 15-bus, 290-bus, and 2224-bus test systems. The maximum testing MSE for the direct physical quantities obtained for all test system sizes is presented in Table II. Consequently, PINN4PF significantly outperforms MLP and LR. Fig. 6 compares the performance of PINN4PF and MLP in terms of direct physical quantities under extreme conditions for the 2224-bus test system. PINN4PF achieves a maximum MSE $[V^{2}]$ 50% lower than MLP for this test system. This improvement is crucial for power system operations, particularly during unexpected events that cause deviations in $v$ from the nominal value. The analysis also considers the performance of PINN4PF, MLP, and LR in approximating derived physical quantities. The maximum testing MSE for all test system sizes is presented in Table III. The table demonstrates the superiority of PINN4PF over MLP and LR across all test system sizes. For the 2224-bus test system, for example, the maximum MSE $[A^{2}]$ obtained by PINN4PF is respectively 92% and 98% lower than MLP and LR. In addition, as the size of the test system increases, the performance gap between PINN4PF, MLP, and LR becomes more pronounced. Fig. 7 compares the performance of PINN4PF and MLP under extreme conditions in the 2224-bus test system for derived physical quantity. It should be noted that while the testing MSE for the direct quantities, i.e., voltage magnitude and phase angle, shows marginal differences between PINN4PF and MLP, PINN4PF significantly outperforms MLP in terms of the derived physical quantities, such as line current.

### v2

@@ IV Proposed architecture: PINN4PF
Figure 1: Schematic diagram of the proposed PINN4PF architecture for PF analysis. The model incorporates two key advancements: (A) a double-head feed-forward NN, $f(\cdot)\in\{f_{0},f_{1},f_{2}\}$, utilizing an adaptive activation function $\sigma^{\prime}(z)=\max(0,\alpha z)$ with a trainable slope $\alpha$; and (B) a physics-informed loss function that combines supervised learning with physical constraints. The input layer receives a concatenated vector of active ($\vec{p}^{d}$) and reactive ($\vec{q}^{d}$) power at all $n$ load buses, resulting in $2n$ input neurons. Shared hidden layers $f_{0}(\cdot)$ act as a feature extractor, projecting the input to a higher-dimensional latent space. Two separate network heads, $f_{1}(\cdot)$ and $f_{2}(\cdot)$, then predict the real ($\vec{\mu}$) and imaginary ($\vec{\omega}$) components of the complex bus voltages, each with $n$ output neurons. The model is trained by minimizing a composite loss that enforces data fidelity (using labels) and physical consistency to enable accurate and generalizable voltage predictions.
@@ IV-A Forward Pass
A double-head feed-forward NN $f(\cdot)\in\{(f_{0},f_{1},f_{2})\}$ is developed to approximate $\hat{\vec{y}}=\{(\hat{\vec{\mu}}_{i},\hat{\vec{\omega}}_{i}):i=1,2,\dots,n\}$ at load buses using the dataset $\{\vec{x},\vec{y}\}$. It has three types of layers, i.e., input, hidden, and output layers, as highlighted in Fig. 1-A. The input and output layers correspond to the input features $\vec{x}$ and output labels $\vec{y}$, respectively. The neurons of the input layer contain active power $\vec{p}^{d}=[p_{1},p_{2},\dots,p_{n}]$ followed by reactive power $\vec{q}^{d}=[q_{1},q_{2},\dots,q_{n}]$ at all load buses, and hence, the input layer has $n\times 2$ neurons, where $n$ is the number of load buses. Following the input layer, there is a set of shared hidden layers $f_{0}(\cdot)$ acting as a feature extractor that projects $\vec{x}$ to a higher-dimensional space, where the two heads, $f_{1}(\cdot)$ and $f_{2}(\cdot)$, separately involve a few hidden layers to respectively approximate $\vec{\mu}=[\mu_{1},\mu_{2},\dots,\mu_{n}]$ and $\vec{\omega}=[\omega_{1},\omega_{2},\dots,\omega_{n}]$. Thus, the output layer of each head has $n$ neurons. The shared set of hidden layers and the two heads each are a chain of functions and can be represented as:
$f_{0}(\vec{x})=l_{m}^{0}\circ\ldots\circ l_{1}^{0}(\vec{x}),$ (3a)
$f_{1}(\vec{x})=l_{m}^{1}\circ\ldots\circ l_{1}^{1}(\vec{x}),$ (3b)
$f_{2}(\vec{x})=l_{m}^{2}\circ\ldots\circ l_{1}^{2}(\vec{x}),$ (3c)
which sequentially process $\vec{x}$ through $m$ hidden layers to obtain $\vec{f}_{0}(\cdot)$, $\vec{f}_{1}(\cdot)$, and $\vec{f}_{2}(\cdot)$, respectively. For the shared set of hidden layers and the two heads, the $k$-th hidden layer ($k=1,2,\dots,m$) is given by:
@@
The Rectified Linear Unit (ReLU) activation function $\sigma(\cdot)$ is applied to the shared set of hidden layers of the two heads (3a). While an adaptive version of ReLU $\sigma^{\prime}(\cdot)$ is developed to apply model-based nonlinear transformations to the two heads (3b) and (3c). That is, a trainable parameter $\alpha$ is introduced to scale the result of the linear transformation while applying nonlinearity. Hence, $\alpha$ serves as a guide to prevent overfitting and improve the generalization ability of the NN against unseen data. $\sigma(\cdot)$ and $\sigma^{\prime}(\cdot)$ are represented as:
$\sigma=\max(0,z),$ (5a)
$\sigma^{\prime}=\max(0,\alpha z),$ (5b)
where $z$ is equivalent to $W_{k}^{T}\cdot\vec{x}+b_{k}$. During each iteration of the training process, the gradients of the loss function concerning the trainable parameters $\nabla_{W_{k},b_{k},\alpha_{k}}L$ are computed for each hidden layer $k$ using the chain rule of differentiation. The resulting gradients are then backpropagated through the network to update and optimize the weights $W_{k}$, biases $b_{k}$, and the scaling parameter $\alpha_{k}$. This process enables the network to learn the optimal values of $\alpha_{k}$, along with $W_{k}$ and $b_{k}$, and hence, potentially leads to improved performance in approximating $\hat{\vec{y}}$. More detailed information about adaptive activation functions can be found in [44, 45].
@@ IV-B Loss Function
The derivation starts with Ohm’s law, which relates voltage $V$, current $I$, and resistance $R$, that is, $V=I\times R$. By extending this equation to the complex domain, $\overline{V}=\overline{I}\times\overline{Z}$ relates the voltage phasor, current phasor, and impedance phasor. From this equation, $\overline{I}$ can be expressed in terms of the voltage phasor and admittance phasor $\overline{Y}$ (inverse of the impedance phasor), that is $\overline{I}=\overline{Y}\times\overline{V}$. The system of equations for all the buses can then be presented in matrix form as:
@@
By rearranging (6), the current flowing into bus $k$ can be presented as a linear combination of the voltages at all other buses with weights given by the corresponding admittance matrix:
@@
Here, the power equation $S=V\times I^{*}$ provides further guides for subsequent derivations, that is $I={S^{*}\over{V^{*}}}$, where $S$ is complex apparent power. The current and voltage phasors are related as:
@@
wherein the right-hand side is an extended form of the right-hand expression in (7). Finally, the expression for calculating the voltage phasor at bus $k$ can be written as:
@@
In (9), $Y_{kk}=G_{kk}+jB_{kk}$ is known from the power system topology. Considering the input features and output labels needed to develop the NN for PF analysis, $V_{k}=\mu_{k}+j\omega_{k}$ and $V_{k}^{*}=\mu_{k}-j\omega_{k}$ are known from the output labels, and $S_{k}^{*}=p_{k}^{d}+jq_{k}^{d}$ is known from the input features. The only unknown expression remaining in (9) is:
@@
which is called hidden function in this study. Note that $\psi$ is unique for each data point. Theoretically, $\psi$ is also unique for each bus. During the training process, the NN first approximates $\psi$ and then uses it to approximate $V=\mu+j\omega$ using:
@@
This unique relation improves the learning process by following the underlying physical laws of the power system. Note that instead of integrating the whole admittance matrix $\begin{bmatrix}Y\end{bmatrix}_{n\times n}$, only the diagonal elements $Y_{kk}$ are needed to obtain the physical loss term, and therefore, the computational resource required is reduced by up to three orders of magnitude compared to when integrating the whole topology information (e.g., [46]). Typically, the goal of the training process is to minimize the difference between $\hat{\vec{y}}$, approximated by the NN, and ground-truth output labels $\vec{y}$, obtained from NR, based on a loss function of choice. Here, a modified loss function is developed that integrates both supervised and physical penalty terms. The former is the mean square of the difference between the output approximated by the NN $f(\cdot)$ and $\vec{y}$. While the latter is the mean square of the difference between the output obtained from the physical model $f^{\prime}(\cdot)$ and $\vec{y}$, as:
@@
where $N$ is the total number of data points $j$. $f(\cdot)\in\{(f_{0},f_{1},f_{2})\}$ and $f^{\prime}(\cdot)$ are the NN architecture (3a)–(3c) and the physical model (11), respectively. $\beta_{0}$ and $\beta_{1}$ represent the coefficients for the supervised and physical penalty terms, respectively, with the constraint that $\beta_{0}+\beta_{1}=1$. Initially, $\beta_{0}=1$ and $\beta_{1}=0$. After 100 epochs, the value of $\beta_{1}$ increases at a specified rate, with its maximum value determined from the sensitivity analysis. $\theta=\{W,b\}$ is the trainable parameters of the NN and $\alpha$ is the trainable parameter for the activation function.
@@ V Results
The performance of PINN4PF is evaluated against two baselines: a linear regression model (LR) and a black-box NN (MLP). Experiments are performed on 4-bus [39], 15-bus [40], 290-bus [41], and 2224-bus [42] test systems. The selected systems contain one reference bus with known voltages in complex number form, $\mu_{0}+j\omega_{0}$, and unknown $p_{0}^{d}$ and $q_{0}^{d}$, and load buses $i$ for which $p_{i}^{d}$ and $q_{i}^{d}$ are known, while complex voltages $\mu_{i}+j\omega_{i}$ are unknown.
Input features and output labels of the datasets are represented as $\vec{x}$ and $\vec{y}$, respectively, where $\vec{x}$ includes $(\vec{p}^{d}_{i},\vec{q}^{d}_{i})$ and $\vec{y}$ includes $(\vec{\mu}_{i},\vec{\omega}_{i})$ obtained from the Newton-Raphson method (NR) for all load buses $i$. The PandaPower Python package [47] is used to perform NR and generate the datasets. Note that PandaPower specifies the state variables of power systems, i.e., $[\delta\text{ }v]^{T}$, that is, there is a need for converting the state variables to $\mu_{i}=v_{i}cos\delta_{i}$ and $\omega_{i}=v_{i}sin\delta_{i}$ to yield the output labels $\vec{y}=\{(\vec{\mu}_{i},\vec{\omega}_{i}):i=1,2,\dots,n\}$.
@@ V-A Model Setup
The training process ends after 5000 epochs for PINN4PF, MLP, and LR. The loss function used is (12), with $\beta_{0}=1$ and $\beta_{1}=0$ for MLP and LR, indicating a supervised penalty term. However, $\beta_{0}$ and $\beta_{1}$ are non-zero for PINN4PF where $\beta_{0}+\beta_{1}=1$, indicating a combination of supervised and physical penalty terms. The activation function is ReLU (5a) for MLP and the shared hidden layers of PINN4PF. The adaptive ReLU (5b) is used for the separated hidden layers of PINN4PF. The Adam optimization algorithm updates the trainable parameters during the training process for PINN4PF, MLP, and LR.
@@ V-B2 Robustness
The robustness of PINN4PF, MLP, and LR is evaluated using the 15-bus test system under varying noise levels. Controlled noise levels ranging from 0% to 10% are introduced to both input features and output labels of the training dataset. The corrupted vectors are defined as $\vec{x}^{\prime}=\vec{x}\pm\vec{r}_{x}$ and $\vec{y}^{\prime}=\vec{y}\pm\vec{r}_{y}$, where $\vec{r}_{x}$ and $\vec{r}_{y}$ are vectors of random values between 0 and 1, and 0 and 0.1, respectively. Fig. 3 illustrates the changes in the MSE for the voltage magnitude $[V^{2}]$ based on the testing dataset with varying noise levels for MLP. The comparison is made relative to a constant curve, representing the MSE obtained for PINN4PF using the highest noise level, i.e., 10%. At the 0% noise level, MLP outperforms PINN4PF trained with 10% noisy data by approximately 18%. However, as the noise level increases, the performance of MLP deteriorates significantly, with MSE increasing up to six times. LR is excluded from the graph as its MSE at different noise levels is two orders of magnitude higher than that of PINN4PF. In addition, LR shows limited improvement with increasing dataset sizes and consistently underperforms compared to both PINN4PF and MLP.
@@ V-B3 Training dataset size
The impact of the size of the training dataset on the performance of MLP is investigated using the 15-bus test system. Fig. 4 illustrates the changes in the MSE for the voltage magnitude $[V^{2}]$ based on the testing dataset with varying training dataset sizes. The comparison is made relative to a constant curve representing the MSE obtained for PINN4PF using 256 training data points. It is observed that MLP requires a training dataset twice as large as that of PINN4PF to achieve a still inferior performance. However, the performance of MLP improves with more training data, reaching a comparable level with a dataset four times larger than that of PINN4PF. This indicates that PINN4PF is more data-efficient. LR is not included in the graph as its MSE with different training dataset sizes is two orders of magnitude larger than that of PINN4PF. In addition, LR shows limited improvement with increasing training dataset sizes and consistently performs worse than PINN4PF and MLP. This indicates that the capacity of LR to capture complex relationships in PF analysis is significantly lower, and increasing the amount of training data is insufficient to bridge the performance gap.
@@ V-B4 Accuracy of derived power flow quantities
For derived physical quantities, i.e., line current, line active power, and line reactive power, the mean and standard deviation of the testing MSE obtained by PINN4PF and MLP are computed and compared in Fig. 5. It is also observed that PINN4PF achieves up to 81%, 63%, and 66% lower maximum testing MSE for the line current $[A^{2}]$, line active power $[W^{2}]$, and line reactive power $[VAR^{2}]$, respectively, compared to MLP.
@@ V-B5 Scalability
The experiments are extended to compare the performance of PINN4PF, MLP, and LR across different test system sizes: 4-bus, 15-bus, 290-bus, and 2224-bus test systems. The maximum testing MSE for the direct physical quantities obtained for all test system sizes is presented in Table II. Accordingly, PINN4PF significantly outperforms MLP and LR. Fig. 6 compares the performance of PINN4PF and MLP in terms of direct physical quantities under extreme conditions for the 2224-bus test system. PINN4PF achieves a maximum MSE $[V^{2}]$ 50% lower than MLP for this test system. This improvement is crucial for power system operations, particularly during unexpected events that cause deviations in $v$ from the nominal value. The analysis also considers the performance of PINN4PF, MLP, and LR in approximating derived physical quantities. The maximum testing MSE for all test system sizes is presented in Table III. The table demonstrates the superiority of PINN4PF over MLP and LR across all test system sizes. For the 2224-bus test system, for example, the maximum MSE $[A^{2}]$ obtained by PINN4PF is respectively 92% and 98% lower than MLP and LR. In addition, as the size of the test system increases, the performance gap between PINN4PF, MLP, and LR becomes more pronounced. Fig. 7 compares the performance of PINN4PF and MLP under extreme conditions in the 2224-bus test system for the derived physical quantity. It should be noted that while the testing MSE for the direct quantities, i.e., voltage magnitude and phase angle, shows marginal differences between PINN4PF and MLP, PINN4PF significantly outperforms MLP in terms of the derived physical quantities, such as line current.